# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 4. Principal part of solutions and perturbed forces
In this section, we construct the principal part of the non-unique steady flow outlined in Section 1.3, and we prove that the perturbative forcing term admits better regularity and smallness than the principal part. The central aim of this section is to establish the following theorem.
@@
For $n\geqslant 3$, there exists a positive constant $\lambda_{*}>0$ such that for every $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant\lambda_{*}$, there exists a divergence-free vector field $V_{\lambda}\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ such that the following (i) and (ii) holds:
@@
    For any $p$ and $q$ satisfying (N1) or (N2) in Theorem 1.2, $V_{\lambda}$ belongs to $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and it holds
@@
    with $\alpha(p):=(3-n/p)(2n+4)-3$ and a positive constant $C=C(n,p,q)$ independent of $\lambda$.
@@
    with $\beta:=6n+10$ and a positive constant $C=C(n,r)$ independent of $\lambda$.
@@
Note that $\alpha(p)\leqslant\alpha(\infty)=6n+9<\beta$, which implies $F_{\lambda}$ is much smaller than $V_{\lambda}$.
@@ 4.1. Notations and key lemmas
Before proving Theorem 4.1, we collect several auxiliary results and notations. We begin by recalling the well-known geometric lemma known as the Nash lemma.
@@
$\mathcal{K}:=\left\{\mathbbm{e}_{i}\ ;\ i=1,\dots,n\right\}$ $\cup\left\{\mathbbm{e}_{i}+\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\}$ (4.5)
$\cup\left\{\mathbbm{e}_{i}-\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\},$ (4.6)
@@
$\Gamma_{k}(M):=\begin{cases}\left(M_{ii}-\dfrac{1}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i},\ i=1,\dots,n),\\[7.0pt] \left(\dfrac{1}{4(n-1)}\pm\dfrac{M_{ij}}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j),\end{cases}$ (4.7)
@@
Here, we introduce some notation that will be used throughout this section.
@@
    Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{R}^{n}\setminus\{0\}$ fulfill $k\cdot a_{k}=0$ and $|a_{k}|\in 2^{\mathbb{N}}$ for all $k\in\mathcal{K}$, and also satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$. For instance, we may choose
    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j).\end{cases}$ (4.8)
@@
    Let $\mu:=2n+4$. For $\lambda\in 2^{\mathbb{N}}$, we define $\{\lambda_{j}\}_{j\in\mathbb{N}_{0}}$ inductively by
@@
    We note that, throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment. We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$. Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in the notation so as to avoid unnecessarily complicating it.
@@
    In what follows, we choose $\lambda$ sufficiently large so that $\max\Lambda_{j}<\min\Lambda_{j+1}$ holds for all $j\in\mathbb{N}_{0}$.
@@
    Note that $h_{j}\sim\log_{2}\log_{\lambda}\lambda_{j}=\lambda_{j-1}^{4\mu}$.
@@
    We choose a $\rho\in C_{c}^{\infty}(\mathbb{R}^{n};[0,\infty))$ so that
@@
    for smooth vector field $v:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Note that $\mathcal{D}v$ is a real symmetric matrix and it holds
@@
In order to define the elements of the desired $V_{\lambda}$, we introduce the following operators.
@@
For $j\in\mathbb{N}$ and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$, we define
${\mathcal{V}}_{j}[v]:=\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\sum_{k\in\mathcal{K}}{\rho}_{j}*\operatorname{div}\mathcal{D}\left({\Gamma}_{j,k}[v]{\Psi}_{j,k}k\right),$ (4.17)
@@
${\Gamma}_{j,k}[v](x):=\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}v(x)\right),$ (4.18)
${\Psi}_{j,k}(x):=\frac{1}{\sqrt{h_{j}}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{\ell}}\psi_{j}(x)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.19)
Let us establish elementary derivative estimates for each $\mathcal{V}_{j}[v]$. Although these bounds are far from optimal, they ensure the well-definedness of Definition 4.6 below, and they provide an essential foundation for obtaining the sharp estimates later on.
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}\leqslant C\lambda_{j}^{2m+3}.$ (4.20)
for all $j\in\mathbb{N}$, $1\leqslant p\leqslant\infty$, and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ with ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$. Moreover, it holds
${\left\|\mathcal{D}{\mathcal{V}}_{j}[v]\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j}^{6}.$ (4.21)
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{-2}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.22)
$\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{2m+2}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.23)
@@
${\left\|\Psi_{j,k}\right\|}_{L^{p}}$ $\leqslant\frac{C}{\sqrt{h_{j}}}\lambda^{\frac{n}{p}j}\sum_{\ell=1}^{\log_{\lambda}\lambda_{j}}\frac{1}{\sqrt{\ell}}$ (4.24)
$\leqslant C\lambda^{-2\mu}_{j-1}\lambda^{\frac{n}{p}j}\sqrt{\log_{\lambda}\lambda_{j}}$ (4.25)
$\leqslant C\lambda^{-2\mu}_{j-1}\lambda_{j},$ (4.26)
@@
${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{5}\leqslant r_{0}\lambda_{j}^{6}$ (4.27)
@@
We now define, by induction, a sequence of flows that will serve as building blocks for the velocity field required in Theorem 4.1.
@@
    Note that (4.21) guarantees that $V_{j-1}$ belongs to the domain of the operator $\mathcal{V}_{j}[\cdot]$, and hence the sequence $\{V_{j}\}_{j\in\mathbb{N}_{0}}$ is well-defined.
@@
    By the definition of $\mathcal{V}_{j}$, we see that each $V_{j}$ is solenoidal.
@@
$V^{(\rm p)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\sum_{k\in\mathcal{K}}\Gamma_{j,k}[V_{j-1}](x)\Psi_{j,k}(x)k$ (4.33)
@@
$V^{(\rm r)}_{j}=V^{({\rm r},1)}_{j}+V^{({\rm r},2)}_{j}$, (4.34)
$\quad V^{({\rm r},1)}_{j}(x):={}\rho_{j}*V^{(\rm p)}_{j}(x)-V^{(\rm p)}_{j}(x),$ (4.35)
$\quad V^{({\rm r},2)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\rho_{j}*\sum_{k\in\mathcal{K}}\operatorname{div}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\right)\cos\left(\lambda_{j}a_{k}\cdot x\right).$ (4.36)
@@
Let $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$ satisfy
@@
$C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\leqslant{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.38)
@@
${\left\|V^{({\rm r})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)$ (4.39)
@@
To begin with, we prepare the $L^{\infty}$ estimates for the derivatives of $\Gamma_{j,k}[V_{j-1}]$. Since $\Gamma_{j,k}[V_{j-1}]$ is constant for $j=1$, we focus on the case of $j\geqslant 2$. Making use of (4.20) for $V_{j-1}=\mathcal{V}_{j-1}[V_{j-2}]$, it holds ${\left\|\nabla^{s}V_{j-1}\right\|}_{L^{\infty}}\leqslant C\lambda_{j-1}^{2s+3}$ ($s\in\mathbb{N}_{0}$). Thus, from the Faà di Bruno formula, it follows that
${\left\|\nabla^{m}\Gamma_{j,k}[V_{j-1}]\right\|}_{L^{\infty}}$ $\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}{\left\|\nabla^{s_{1}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}\cdots{\left\|\nabla^{s_{\ell}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}$ (4.41)
$\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}\cdot{\lambda_{j-1}^{2(s_{1}+1)+3}}\cdots{\lambda_{j-1}^{2(s_{\ell}+1)+3}}$ (4.42)
$=C\lambda_{j-1}^{2m}\sum_{\ell=1}^{m}\lambda_{j-1}^{-\ell}$ (4.43)
@@
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.45)
$\quad+C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.46)
$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.47)
$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.48)
@@
${\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}\leqslant C\lambda_{j-1}^{4n}.$ (4.50)
@@
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\max_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.51)
$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-1}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.52)
$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.53)
$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1-C^{2}\lambda_{j-1}^{2-\mu}-C^{2}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.54)
$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.55)
@@
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s+1,q)$ (4.60)
$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.61)

### v2

@@ 4. Principal part of solutions and perturbed forces
In this section, we construct the principal part of the non-unique steady flow outlined in Section 1.3, and prove that the perturbative forcing term admits better regularity and smallness than the principal part. The central aim of this section is to show the following theorem.
@@
For $n\geqslant 3$, there exists a constant $\lambda_{*}>0$ such that for every $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant\lambda_{*}$, there exists a divergence-free vector field $V_{\lambda}\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ such that the following (i) and (ii) hold true.
@@
    For any $(p,q)$ satisfying (N1) or (N2) in Theorem 1.2, $V_{\lambda}$ belongs to $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and it holds
@@
    with $\alpha(p):=(3-n/p)(4n+5)-4-n/p$ and some constant $C=C(n,p,q)>0$.
@@
    with $\beta:=12n+12$ and some constant $C=C(n,r)>0$.
@@
Since $\alpha(p)\leqslant\alpha(\infty)=12n+11<\beta$ and $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})\hookrightarrow\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$, $F_{\lambda}$ is much smaller and more regular than $V_{\lambda}$.
@@ 4.1. Notations and key lemmas
Before proving Theorem 4.1, we collect several auxiliary results and notations. We begin by recalling the well-known geometric lemma.
@@
$\mathcal{K}:=\left\{\mathbbm{e}_{i}\ ;\ i=1,\dots,n\right\}$ $\cup\left\{\mathbbm{e}_{i}+\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i<j\right\}$ (4.5)
$\cup\left\{\mathbbm{e}_{i}-\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i<j\right\},$ (4.6)
@@
$\Gamma_{k}(M):=\begin{cases}\left(M_{ii}-\dfrac{1}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i},\ i=1,\dots,n),\\[7.0pt] \left(\dfrac{1}{4(n-1)}\pm\dfrac{M_{ij}}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j),\end{cases}$ (4.7)
@@
Here, we introduce some notations that are to be used throughout this section.
@@
    Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{S}^{n-1}$ satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$. For instance, we may choose
    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
@@
    Let $\mu=4n+5$. For $\lambda\in 2^{\mathbb{N}}$, we define $\{\lambda_{j}\}_{j\in\mathbb{N}_{0}}$ inductively by
@@
    We note that, throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment. We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$. Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in their notations so as to avoid unnecessarily complicating them.
@@
    In what follows, we choose $\lambda$ sufficiently large so that $\Lambda_{j}\neq\varnothing$ for all $j\in\mathbb{N}$.
@@
    Note that there holds
    $h_{j}\sim{}$ $\log_{2}\frac{\log_{\lambda}\lambda_{j}}{\mu\log_{\lambda}\lambda_{j-1}}$ (4.12)
    $={}$ $\begin{cases}\lambda_{j-1}^{4\mu}-\log_{2}\mu&(j=1),\\ \lambda_{j-1}^{4\mu}-\lambda_{j-2}^{4\mu}-\log_{2}\mu,\quad&(j=2,3,4,\dots),\end{cases}$ (4.13)
    which yields $h_{j}\sim\lambda_{j-1}^{4\mu}$.
@@
    We choose a function $\rho\in C_{c}^{\infty}(\mathbb{R}^{n};[0,\infty))$ so that
@@
    for all smooth vector field $v:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Note that $\mathcal{D}v$ is a real symmetric matrix and it holds
@@
In order to define the building blocks of the desired $V_{\lambda}$, we introduce the following operators.
@@
For $j\in\mathbb{N}$ and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{8}$, we define
${\mathcal{V}}_{j}[v]:=\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\operatorname{div}\mathcal{D}\left({\Gamma}_{j,k}[v]{\Psi}_{j,k}k\right),$ (4.19)
@@
${\Gamma}_{j,k}[v](x):=\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}v(x)\right),$ (4.20)
${\Psi}_{j,k}(x):=\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\psi_{j}(x)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.21)
Remark 4.5.
When $\mathbb{R}^{n}$ is replaced by $\mathbb{T}^{n}$, it is not necessary to multiply by $\psi_{j}$ in the definition of $\Psi_{j,k}$.
Let us establish elementary derivative estimates for each $\mathcal{V}_{j}[v]$, which rely on the mollifier. Although these bounds are far from optimal, they ensure the well-definedness of Definition 4.7 below, and they provide an essential foundation for obtaining the sharp estimates later on.
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}\leqslant C\lambda_{j}^{2m+4}.$ (4.22)
for all $j\in\mathbb{N}$, $1\leqslant p\leqslant\infty$, and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ with ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{8}$. Moreover, it holds
${\left\|\mathcal{D}{\mathcal{V}}_{j}[v]\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j}^{8}.$ (4.23)
@@
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
$\leqslant C\lambda_{j-1}^{4}\lambda_{j}^{2m+4}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.25)
@@
${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{-4\mu}_{j-1}$ (4.26)
@@
${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
@@
We now define the building blocks for the velocity field required in Theorem 4.1.
@@
    Note that (4.23) guarantees that $V_{j-1}$ belongs to the domain of the operator $\mathcal{V}_{j}$, and hence the sequence $\{V_{j}\}_{j\in\mathbb{N}_{0}}$ is well-defined.
@@
    By $\operatorname{div}\mathcal{D}=-\Delta\mathbb{P}$ in the definition of $\mathcal{V}_{j}$, we see that each $V_{j}$ is solenoidal.
@@
$V^{(\rm p)}_{j}(x):={}$ $\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)\operatorname{div}\mathcal{D}\left(k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)\right)$ (4.33)
$={}$ $\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)$ (4.34)
@@
$V^{({\rm r},1)}_{j}(x):={}\rho_{j}*V^{(\rm p)}_{j}(x)-V^{(\rm p)}_{j}(x),$ (4.35)
$\begin{aligned} V^{({\rm r},2)}_{j}(x):={}&\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{\ell}\sqrt{h_{j}\ell}}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)\nabla\cos\left(\lambda^{\ell}a_{k}\cdot x\right)\\ ={}&-\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{\ell}\sqrt{h_{j}\ell}}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)a_{k}\sin\left(\lambda^{\ell}a_{k}\cdot x\right),\end{aligned}$ (4.36)
$V^{({\rm r},3)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\operatorname{div}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.37)
We also introduce a function
$\Phi_{j,k,\ell}(x):=\frac{1}{\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)$ (4.38)
so that $V^{({\rm p})}_{j}$ has the following simplified representation:
$V^{(\rm p)}_{j}(x)=\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\Phi_{j,k,\ell}(x).$ (4.39)
@@
Let $1\leqslant p,q\leqslant\infty$ and let $s\in\mathbb{R}$ satisfy
@@
$C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\leqslant{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q),$ (4.41)
@@
${\left\|V^{({\rm r})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)$ (4.42)
@@
To begin with, we prepare the $L^{\infty}$ estimates for the derivatives of $\Gamma_{j,k}[V_{j-1}]$. Since $\Gamma_{j,k}[V_{j-1}]$ is constant for $j=1$, we focus on the case of $j\geqslant 2$. Making use of (4.22) for $V_{j-1}=\mathcal{V}_{j-1}[V_{j-2}]$, it holds ${\left\|\nabla^{s}V_{j-1}\right\|}_{L^{\infty}}\leqslant C\lambda_{j-1}^{2s+4}$ ($s\in\mathbb{N}_{0}$). Thus, from the Faà di Bruno formula, it follows that
${\left\|\nabla^{m}\Gamma_{j,k}[V_{j-1}]\right\|}_{L^{\infty}}$ $\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-8\ell}{\left\|\nabla^{s_{1}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}\cdots{\left\|\nabla^{s_{\ell}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}$ (4.44)
$\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-8\ell}\cdot{\lambda_{j-1}^{2(s_{1}+1)+4}}\cdots{\lambda_{j-1}^{2(s_{\ell}+1)+4}}$ (4.45)
@@
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.47)
$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.48)
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\left(1+\lambda^{n(1-\frac{n}{p})j}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.49)
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q),$ (4.50)
@@
${\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}\leqslant C\lambda^{nj}\lambda_{j-1}^{4n}.$ (4.52)
@@
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\geqslant C^{-1}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\max_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.53)
$\quad-C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.54)
$\quad-C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.55)
$\geqslant C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\left(1-C^{2}\lambda^{n(1-\frac{1}{p})j}\lambda_{j-1}^{2-\mu}-C^{2}\lambda^{n(1-\frac{1}{p})j}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.56)
$\geqslant C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q).$ (4.57)
@@
$\leqslant C\lambda_{j}^{-2}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s+1,q)$ (4.62)
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-1,q).$ (4.63)